Mrs. Jones is 3 times as old as her son. Ten years ago she was 5 times as old as her son was then. Find each of their ages.
Answer by Alan3354(69443)

Mrs. Jones is 3 times as old as her son. Ten years ago she was 5 times as old as her son was then. Find each of their ages.
------------
j = her age
s = son's age
---
j = 3s
j-10 = 5*(s-10)
Sub for j
3s-10 = 5s-50
etc